OpenStudy (ivonne123):

Find the least common multiple of the polynomials. please 6y^2 , 21y

OpenStudy (mathmale):

What do 6 and 21 have in common? What do y^2 and y have in common?

OpenStudy (jdoe0001):

hint: \(\bf \begin{cases} 6y^2\to {\color{blue}{ 3}}\cdot 2\cdot {\color{blue}{ y}}\cdot y \\ \quad \\ \quad \\ 21y\to {\color{blue}{ 3}}\cdot 7\cdot {\color{blue}{ y}} \end{cases}\)

OpenStudy (mathmale):

Again: What do 6 and 21 have in common? What do y^2 and y have in common?